# RadZero: Similarity-Based Cross-Attention for Explainable Vision-Language Alignment in Radiology with Zero-Shot Multi-Task Capability

### v1

@@ 6 Conclusion
In this work, we introduced RadZero, a novel similarity-based cross-attention framework for vision-language alignment in radiology. By computing cosine similarities between text descriptions and local image patches, RadZero enhanced interpretability while demonstrating remarkable zero-shot capability across classification, grounding, and segmentation. Its ability to process high-resolution chest X-rays combined with a multi-positive contrastive learning strategy, enabled effective representation learning without requiring pixel-level annotations. Extensive evaluations on public chest radiograph benchmarks revealed RadZero’s superiority over state-of-the-art methods in zero-shot classification, grounding, and segmentation. Furthermore, cross-modal similarity map analysis highlighted its advantage in explainability, as the similarity map provide a transparent rationale for how conclusions are derived. Qualitative assessments further demonstrated RadZero’s potential for open-vocabulary semantic segmentation, validating its adaptability to diverse medical imaging scenarios.
@@
RadZero, while achieving impressive results, has limitations that indicate areas for future research. It did not consistently surpass all benchmarks in zero-shot classification, emphasizing the need for improved generalization. Fine-tuning with class labels could facilitate the development of a high-performance explainable classifier. Extending the similarity-based cross-attention approach to imaging modalities like CT or MRI offers potential for interpretable vision-language models in medical imaging.
Supplementary Material
Appendix A Abbreviation
@@
Appendix B Model Training Details
Appendix C Ablation Studies
Impact of multi-positive pairs for contrastive learning.
To apply contrastive learning, CARZero [19] randomly sampled one sentence from multiple available options. We compare this random selection strategy with our multi-positive NCE loss, which leverages all sentences associated with an image. As shown in Table 6, utilizing multiple sentences consistently outperformed random selection across all tasks, demonstrating the effectiveness of our approach.
  Method class. ground. phrase. seg.
Random Select 0.827 0.586 0.832 0.138
Multi-positive 0.841 0.622 0.844 0.171
Table 6: Impact of multi-positive pairs for contrastive learning.
Impact of view position.
  View Position class. ground. phrase. seg.
Frontal 0.831 0.604 0.838 0.161
All View 0.841 0.622 0.844 0.171
Table 7: Impact of view position.
Impact of image encoder and resolution.
DINOv2-based encoders significantly outperformed M3AE across all tasks, demonstrating the effectiveness of the DINOv2 pretraining strategy for chest X-ray representation learning. Comparing XrayDINOv2 at resolutions of 224 and 518, we observe that higher image resolution improves fine-grained tasks such as grounding and segmentation. RadDINO and XrayDINOv2 showed similar performance, suggesting that our approach is effectively applied to models trained with the DINOv2 strategy on chest X-ray images.
  Image Image class. ground. phrase. seg.
Encoder Resolution
M3AE [11] 224 0.703 0.284 0.647 0.027
RadDINO [26] 518 0.850 0.610 0.844 0.144
XrayDINOv2 [5] 224 0.841 0.548 0.832 0.118
XrayDINOv2 518 0.841 0.622 0.844 0.171
Table 8: Impact of image encoder and resolution.
Impact of trainable vision layer architecture.
  Model class. ground. phrase. seg.
Linear 0.826 0.549 0.826 0.100
1 Transformer layer 0.835 0.585 0.832 0.158
2 Transformer layer 0.841 0.622 0.844 0.171
Table 9: Impact of trainable vision layer architecture.
Impact of text encoder.
  Image class. ground. phrase. seg.
Encoder
BioBERT 0.842 0.582 0.832 0.127
MPNet 0.841 0.622 0.844 0.171
Table 10: Impact of text encoder.
Impact of batch size.
  Batch size class. ground. phrase. seg.
64 0.835 0.583 0.826 0.165
128 0.840 0.594 0.850 0.177
256 0.841 0.622 0.844 0.171
Table 11: Impact of batch size.
@@ Appendix C Ablation Studies
Appendix D Prompt for finding-sentence extraction.
Appendix E Additional Visualization Results
Figure 9 presents cross-modal similarity maps for 10 different findings, following the pipeline in Sec. 5.2. The highest similarity regions align well with the bounding boxes, even for multiple or small lesions. The similarity probability was above 0.5 for all findings except calcification. While the model correctly localized calcifications, the activated regions appeared as small bright spots, leading to a lower similarity probability of 0.45 due to the weighted sum calculation. This highlights a limitation of RadZero, suggesting the need for further refinement in future work.
@@ Appendix E Additional Visualization Results
Figure 9: Cross-modal similarity maps for 10 findings. Visualization of similarity maps generated by RadZero on the ChestXDet10 dataset. Red boxes indicate ground truth bounding boxes. The similarity probability $\hat{l}$ is shown in the top-right corner of each map.

### v2

@@ 6 Conclusion
In this work, we introduced RadZero, a novel VL alignment model for radiology that achieved strong zero-shot performance in classification, grounding, and segmentation. Central to RadZero is VL-CABS, which computes image-text similarity at the patch-level to improve interpretability. Combined with multi-positive contrastive training, VL-CABS enabled effective representation learning without pixel-level annotations, and the support for high-resolution inputs further boosted performance. Extensive evaluations on public chest radiograph benchmarks showed that RadZero outperformed SOTA methods. VL similarity map analysis highlighted the enhanced explainability of VL-CABS by providing transparent rationales for how conclusions are derived. Qualitative assessments further demonstrated RadZero’s potential for open-vocabulary semantic segmentation.
@@
Despite its impressive results, RadZero has limitations that indicate areas for future research. The performance lag on certain datasets highlights the need for improved generalization. Its reliance on the pre-trained vision encoder is also a constraint. Fine-tuning with class labels could facilitate the development of a high-performance explainable classifier. Extending VL-CABS to general image domains offers potential for more interpretable VLMs across diverse visual tasks.
Appendix A Are Attention Maps of Medical VLMs Explainable?
Explainability is critical for clinical deployment of deep learning models, and medical VLMs have adopted attention maps as their de facto interpretable features [36, 19, 41]. However, attention maps alone reveal only where the model is focusing, not why it makes a particular prediction. Without the underlying image–text similarity scores, such heatmaps lack interpretability. Interpreting attention maps for complex text queries can be unintuitive and may depend on access to ground-truth labels.
Figure 6: Attention maps from CARZero [19] Image–text similarity probabilities are obtained by applying a sigmoid function to the classification logits.
To illustrate the challenges of interpreting attention maps, Figure 6 visualizes the attention outputs of CARZero [19] for different image–text pairs, with the overlaid values indicating the corresponding similarity probabilities. In (a), the prompt “There is right effusion” produces a heatmap focused on the right lower lung and a high similarity probability (0.560), intuitively linking the attended region to the predicted finding. In (b), when queried with “There is left effusion” on the same image, the attention map still highlights the right lower lung. While we can observe where the model attends, the attention map alone offers no clear explanation for why that region is relevant. The image–text similarity probability for the prompt “There is left effusion” is low (0.048), which—when considered alongside the attention map—can be interpreted as indicating that the attended right lung region does not support the presence of left-sided effusion. However, how should we interpret the fact that the model attends to the right lung when queried about left effusion? Making sense of this behavior requires access to the ground truth: effusion is present in the right lower lung but absent on the left. With this context, one might infer that the model correctly identifies effusion on the right, recognizes that it does not match the left-sided query, and therefore assigns a low similarity score.
This example underscores the need to interpret attention maps in conjunction with image–text similarity scores. For complex query texts, meaningful interpretation often depends on knowing the ground truth—a significant limitation in medical VLMs where such labels are frequently unavailable. This reliance undermines the standalone explainability of attention visualizations. To address these limitations, we propose Vision–Language Cross-Attention Based on Similarity (VL-CABS), a framework that enables transparent inspection of the vision–language decision process. Detailed examples and analysis of its application in RadZero are provided in Section 5.2.
Appendix B Model Training Details
Appendix C Ablation Studies
Impact of view position.
Impact of image encoder and resolution.
DINOv2-based encoders significantly outperformed M3AE across all tasks, demonstrating the effectiveness of the DINOv2 pre-training strategy for chest X-ray representation learning. Comparing XrayDINOv2 at resolutions of 224 and 518, we observe that higher image resolution improves fine-grained tasks such as grounding and segmentation. RadDINO and XrayDINOv2 showed similar performance, suggesting that our approach is effectively applied to models trained with the DINOv2 strategy on chest X-ray images.
Impact of trainable vision layer architecture.
  View Position class. ground. phrase. seg.
Frontal 0.831 0.604 0.838 0.161
All View 0.841 0.622 0.844 0.171
Table 6: Impact of view position.
  Image Image class. ground. phrase. seg.
Encoder Resolution
M3AE [11] 224 0.703 0.284 0.647 0.027
RadDINO [27] 518 0.850 0.610 0.844 0.144
XrayDINOv2 [5] 224 0.841 0.548 0.832 0.118
XrayDINOv2 518 0.841 0.622 0.844 0.171
Table 7: Impact of image encoder and resolution.
  Model class. ground. phrase. seg.
Linear 0.826 0.549 0.826 0.100
1 Transformer layer 0.835 0.585 0.832 0.158
2 Transformer layers 0.841 0.622 0.844 0.171
Table 8: Impact of trainable vision layer architecture.
  Image class. ground. phrase. seg.
Encoder
BioBERT 0.842 0.582 0.832 0.127
MPNet 0.841 0.622 0.844 0.171
Table 9: Impact of text encoder.
  Batch size class. ground. phrase. seg.
64 0.835 0.583 0.826 0.165
128 0.840 0.594 0.850 0.177
256 0.841 0.622 0.844 0.171
Table 10: Impact of batch size.
Impact of text encoder.
Impact of batch size.
Appendix D Prompt for finding-sentence extraction.
@@ Appendix C Ablation Studies
Appendix E Abbreviation
Table 11 lists the abbreviations used for lesions and anatomical regions. The left column shows the abbreviated terms, and the right column gives their description.
Appendix F Additional Visualization Results
@@
Figure 10 presents VL similarity maps for 10 different findings, following the pipeline in Sec. 5.2. The highest similarity regions align well with the bounding boxes, even for multiple or small lesions. The similarity probability was above 0.5 for all findings except calcification. While the model correctly localized calcifications, the activated regions appeared as small bright spots, leading to a lower similarity probability of 0.45 due to the weighted sum calculation. This highlights a limitation of RadZero, suggesting the need for further refinement in future work.
@@
Figure 10: VL similarity maps for 10 findings. Visualization of similarity maps generated by RadZero on the ChestXDet10 dataset. Red boxes indicate ground truth bounding boxes. The similarity probability $\hat{l}$ is shown in the top-right corner of each map.